# Simple Lifelong Learning Machines

### v21

@@ A-E Real Data Extended Experiments and Details
CIFAR 10X10 CIFAR 100 $5000$ $10000$ $3\times 32\times 32$
5-dataset CIFAR-10 $50000$ $10000$ $3\times 32\times 32$ (resized)
MNIST $60000$ $10000$
SVHN $73257$ $26032$
notMNSIT $16853$ $1873$
Fashion-MNIST $60000$ $10000$
Split Mini-Imagenet Mini-Imagenet $48000$ $12000$ $3\times 84\times 84$
FOOD1k 50X20 Food1k $60000$ $99682$ $3\times 50\times 50$ (resized)
Spoken Digit Spoken Digit $1650$ $1350$ $28\times 28$ (processed and resized)
@@
In this experiment, we have used the Mini-Imagenet dataset [65]. The dataset was split into $20$ tasks with $5$ classes each. Each task has $2400$ training samples and $600$ testing samples. As shown in Figure 5 right column, we get positive forward and backward transfer for SiLLy-N. However, although samples per task is lower compared to that of 5-dataset, it is still quite high. Hence, Model Zoo outperforms all the algorithms in this experiment.
@@
In this experiment, we have used 5-dataset [65]. It consists of $5$ tasks from five different datasets: CIFAR-10 [67], MNIST, SVHN [76], notMNIST [77], Fashion-MNIST [78]. All the monochromatic images were converted to RGB format, and then resized to $3\times 32\times 32$. As shown in Appendix Table III, training samples per task in $5$-dataset is relatively higher than that of low data regime typically considered in lifelong learning setting. However, as shown in Figure 5 left column, SiLLy-N show less forgetting than most of the reference algorithms. On the other hand, Model Zoo shows comparatively better performance in relatively high task data size setup. Recall that SiLLy-N is based on bagging, and Model Zoo is based on boosting. It is well known that boosting often outperforms bagging when sample sizes are large  [^2] Authors in [79] shows that both bagging and boosting asymptotically converge to the Bayes optimal solution. However, for finite sample size and similar model complexity, we empirically find bagging approach to lifelong learning performs better than that of boosting when the training sample size is low whereas boosting performs better on large training sample size (See Figure 1). This is consistent with similar results in single task learning [80, 81, 41].
@@
In this experiment, we used the Spoken Digit dataset provided in https://github.com/Jakobovski/free-spoken-digit-dataset. The dataset contains audio recordings from six different speakers with $50$ recordings for each digit per speaker ($3000$ recordings in total). The experiment was set up with six tasks where each task contains recordings from only one speaker. For each recording, a spectrogram was extracted using Hanning windows of duration $16$ ms with an overlap of $4$ ms between the adjacent windows. The spectrograms were resized down to $28\times 28$. The extracted spectrograms from eight random recordings of ‘$5$’ for six speakers are shown in Figure 6. For each Monte Carlo repetition of the experiment, spectrograms extracted for each task were randomly divided into $55\%$ train and $45\%$ test set. The experiment is summarized in Figure 7. Note that we could not run the experiment on other $5$ reference algorithms using the code provided by their authors.

### v22

@@ A-E Real Data Extended Experiments and Details
CIFAR 10X10 CIFAR 100 500050005000 100001000010000 3×32×323$\times$ 32$\times$ 323 × 32 × 32
5-dataset CIFAR-10 500005000050000 100001000010000 3×32×323$\times$ 32$\times$ 323 × 32 × 32 (resized)
MNIST 600006000060000 100001000010000
SVHN 732577325773257 260322603226032
notMNSIT 168531685316853 187318731873
Fashion-MNIST 600006000060000 100001000010000
Split Mini-Imagenet Mini-Imagenet 480004800048000 120001200012000 3×84×843$\times$ 84$\times$ 843 × 84 × 84
FOOD1k 50X20 Food1k 600006000060000 996829968299682 3×50×503$\times$ 50$\times$ 503 × 50 × 50 (resized)
Spoken Digit Spoken Digit 165016501650 135013501350 $\times 28$ 28 (processed and resized)
@@
In this experiment, we have used the Mini-Imagenet dataset [67]. The dataset was split into 202020 tasks with 555 classes each. Each task has 240024002400 training samples and 600600600 testing samples. As shown in Figure 5 right column, we get positive forward and backward transfer for SiLLy-N. However, although samples per task is lower compared to that of 5-dataset, it is still quite high. Hence, Model Zoo outperforms all the algorithms in this experiment.
@@
In this experiment, we have used 5-dataset [67]. It consists of 555 tasks from five different datasets: CIFAR-10 [69], MNIST, SVHN [78], notMNIST [79], Fashion-MNIST [80]. All the monochromatic images were converted to RGB format, and then resized to 3×32×323$\times$ 32$\times$ 323 × 32 × 32. As shown in Appendix Table III, training samples per task in 555-dataset is relatively higher than that of low data regime typically considered in lifelong learning setting. However, as shown in Figure 5 left column, SiLLy-N show less forgetting than most of the reference algorithms. On the other hand, Model Zoo shows comparatively better performance in relatively high task data size setup. Recall that SiLLy-N is based on bagging, and Model Zoo is based on boosting. It is well known that boosting often outperforms bagging when sample sizes are large  [^2] Authors in [81] shows that both bagging and boosting asymptotically converge to the Bayes optimal solution. However, for finite sample size and similar model complexity, we empirically find bagging approach to lifelong learning performs better than that of boosting when the training sample size is low whereas boosting performs better on large training sample size (See Figure 1). This is consistent with similar results in single task learning [82, 83, 43].
@@
[…]